# Towards more realistic evaluation of LLM-based code generation: an experimental study and beyond

### v1

@@ 3.2.1. Projects selection
In order to make HumanEvo more representative and adequate, we adopt a comprehensive approach to selecting functions from various open-source projects. For Python, we select high-quality real-world projects from a curated list of the top 5,000 most downloaded PyPI libraries during 2023111https://hugovk.github.io/top-pypi-packages/. Subsequently, we deploy the selected projects and run their testing frameworks to ensure each of them can run correctly and stably. We then proceed to collect PRs for these repositories through the GitHub developer API [^2] https://api.github.com. For Java, as widely recognized projects typically indicate extensive documentation, structured open-source development guidelines, and functional, well-formatted code, our approach involves searching for the top 200 starred projects on GitHub. Then, we deploy every project on the list to select projects that can be successfully deployed and pass the corresponding test suite. This methodology ensures that the functions selected for evaluation in HumanEvo are representative of real-world programming scenarios, encompassing a diverse range of projects and languages.
3.2.2. PR crawling
@@
We crawl a large number of initial PRs from the popular open-source repositories selected in the procedure above. Specifically, we obtain 12,039 PRs and meticulously record metadata pertaining to these PRs for subsequent filtering processes.
3.2.3. PR attribute-based filtering
@@
In this step, we preliminarily filter the crawled PRs based on the following three attributes.
@@
  • The PR must introduce new functions. These newly added functions are typically intended to develop the software project by adding new features, improvements, or fulfilling general development objectives, which can be selected as programming questions.
@@
3.2.4. PR execution-based filtering
In this step, we perform execution-based filtering for the PRs filtered in the attributes filtering step.
Firstly, we prepare the execution environment for each candidate task instance in HumanEvo. For each task instance, we first determine the version of the project corresponding to the PR, then manually identify the required dependency environment for running the project at that version. For example, for a Python task instance, we confirm the required Python version and the types and versions of third-party libraries it depends on. After obtaining this dependency environment information, we create a virtual execution environment for each version of the project. Task instances that can not install the dependency environment properly are filtered out during this step. After creating the execution environment, we execute the project’s testing framework in the corresponding environment.
@@ 3.2.4. PR execution-based filtering
3.2.5. Function selection
In this step, we manually select functions from the PRs that have passed the filtering process above to serve as programming questions. After completing the above filtering, we confirm that the remaining PRs meet our requirements. Therefore, we select newly added functions from these PRs to as programming questions. During the selection process, to avoid choosing basic functions, we filter out initialization functions, constructors, and destructors, to ensure the quality of the selected programming questions. Finally, we acquire a total of 400 programming questions.
3.2.6. Human annotation
In this step, we recruit six annotators with more than five years of programming experience to manually rewrite two different styles of docstrings for the selected functions and label each function with a dependency level.
We rewrite the docstring for each instance in HumanEvo to ensure the quality of the benchmark dataset and mitigate data leakage issues as much as possible. Specifically, in the process of data collection, we find that the docstring styles of actual projects can be divided into two categories: the detail and the brief, so in order to make our benchmark more diverse and better reflect the actual development scenarios, we provide two types of the docstring for each task instance. The detailed docstring describes the function of the target function in detail, as well as the types and meanings of its inputs and outputs. In contrast, a brief docstring only briefly describes what the target function does. After the annotators complete writing the docstrings, we conduct a cross-validation review, requiring each task instance to be verified and confirmed by two developers to ensure accuracy. If the annotators disagree with a docstring written by the previous annotator during the review process, they will discuss the issue together until all annotators reach an agreement.
@@ 3.2.6. Human annotation
3.3. Benchmark Characteristics
3.3.1. Evolving-aware benchmark
@@
HumanEvo is the pioneering repository-level code generation benchmark to introduce the evolving-aware concept. We highlight a significant oversight in prior research: the developmental trajectory of a project is an evolving process, altering the project context available to programmers as the project advances through various stages. Previous studies have overlooked this temporal dimension, resulting in LLMs being fed with code that had not yet been implemented in the project when certain functions were developed. Consequently, it also misses some context that should have been present but might have been deleted or updated during the project’s evolution. This undoubtedly leads to the model facing an unrealistic development scenario. As a result, the performance of LLMs in real-world software development tasks remains unknown. Therefore, we are committed to enhancing HumanEvo to serve as a benchmark that better simulates real-world software code development scenarios, aiming to more accurately reflect the true performance of LLMs applied to pragmatic development tasks.
3.3.2. Strict data filtering process
@@
We ensure the quality of the selected function from several aspects. First, we ensure that the selected projects are of high quality because high-quality projects generally have excellent maintenance, comprehensive testing frameworks, and good coding styles. For Python, we choose the projects with the highest download count in PyPI. For Java, we select projects from the top 200 projects on GitHub based on the number of stars, with an average of around 24k stars. Previous benchmarks have ensured the functionality of target functions by writing unit tests for each function separately. However, we believe that in a project with complex dependencies, adding a new function may have certain impacts on other parts of the project. Relying solely on unit tests for the target function is insufficient to ensure that the project can still function properly and pass the project’s test framework after inserting the function. Therefore, in this work, we directly run the project’s test framework to ensure the correctness of the target function’s functionality while ensuring that inserting the target function does not have a negative impact on other parts of the project. Of course, to ensure that the test framework covers the target function, during the function selection process, we verify that our selected functions are covered by the project’s test framework through execution-based validation. In other words, a “fail-to-pass” test result implies that the test is set up for the target function and the target function has passed the test.
Table 1. Dependency level in HumanEvo.
@@
Standalone 46 (23.0%) 26 (13.0%)
Intra-class 61 (30.5%) 61 (30.5%)
Inter-class 103 (51.5%) 113 (56.5%)
3.3.3. Human annotation
@@
Furthermore, to facilitate the assessment of LLMs’ capability in generating functions with varying levels of dependency, we categorize the selected functions based on their dependency levels, including standalone, intra-class, and inter-class. Standalone functions can be implemented without relying on other functions within the project. Intra-class functions require dependencies on other functions within the same class. Inter-class functions imply more complex dependencies, with most of the dependent functions being located in other files within the project. Manual categorization of dependency levels aids in analyzing the performance of LLMs in generating functions of varying complexity, facilitating researchers in improving the application of LLMs in practical development scenarios based on different contexts. Moreover, as indicated in Table 1, the majority of our benchmark consists of functions with deep dependency levels. This suggests that our benchmark presents a sufficiently challenging scenario.
4. Experimental Design
@@
4.1. Model Selection
We categorize mainstream LLMs for code generation into two groups: open-source LLMs and closed-source LLMs. For open-source LLMs, we select the CodeLlama series (including CodeLlama-7B, CodeLlama-13B, and CodeLlama-34B) and the DeepSeekCoder series (including DeepSeekCoder-6.7B and DeepSeekCoder-33B). The CodeLlama and DeepSeekCoder series are popular models that perform well in code generation. For closed-source LLMs, we choose the commonly used commercial models: GPT-3.5-Turbo and GPT-4, with GPT-4 having demonstrated superior performance levels in previous benchmarks.
4.2. Research Questions
Our experiments intend to answer the following research questions:
@@ 4.2. Research Questions
    We conduct experiments under two different settings, evolving-aware and evolving-ignored, to explore the differences in model performance between these two settings and reveal the actual performance of the models in real development scenarios.
@@
    We selected multiple versions of the project after the target function was committed as the context source to explore how the model’s performance gradually deviates from reality over time under the evolving-ignored setting.
@@
4.3. Experimental Settings
Due to the large size of modern software repositories, it is not feasible to feed the entire repository to the model. Thus, previous researchers mainly obtain project context that might be beneficial for generating target functions through two approaches: retrieval-based and static analysis-based methods. In the experiments below, we cover both of these technical routes to demonstrate that the evolving-ignored situation leads to the deviation of the context source from reality.
4.3.1. Retrieval based project context acquisition
We adopt the standard retrieval-augmented-generation pipeline for repository-level code generation (Zhang et al., 2023b; Formal et al., 2022; Santhanam et al., 2022; Xiong et al., 2021). In the evolving-aware setting, we restore the project repository to the state before the target code was committed for each task instance. On the contrary, in the evolving-ignored settings, we directly utilize the latest project version as the context source. Then, we construct a corresponding context retrieval database by partitioning the source code files from the repository into a collection of code snippets. Subsequently, we utilize the docstring of the target code as the query for retrieval and compute the Jaccard similarity between the query and all code snippets in the context retrieval database. After sorting all the code snippets by similarity score, we select project contexts with higher similarity scores as project context for target function generation.
4.3.2. Static analysis based project context acquisition
Similar to the static analysis based repository-level code generation work (Shrivastava et al., 2023), we obtain code files in the repository that might help generate the target code through static analysis. First, we roll back the project repository to the corresponding state according to evolving-ignored or evolving-aware settings. Then, we extract the source code from these files as project context to feed to the model. There are four types of context sources: local file (the file where the target code resides), import file (files imported by the local file in the project repository), sibling file (files located in the same folder as the local file), and similar file (files in the project repository with names similar to the local file). We utilize four prompting strategies from (Shrivastava et al., 2023): local – extracting context solely from the local file; local+import – extracting code from both local and import files; local+sibling - extracting code from both local and sibling files; local+sim - extracting code from both local and similar files.
4.3.3. Evaluation metric
We assess the functionality correctness of code generated by LLMs with execution-based evaluation and utilize the commonly used metric Pass@1 for evaluation. Note that, to mitigate issues stemming from model-generated randomness, the experimental results presented subsequently are obtained by conducting three repeated experiments and averaging the results.
Table 2. Retrieval-based project context acquisition approach: model performance under evolving-aware (HumanEvo) and evolving-ignored (EI) settings. DepSCode is short for DeepSeekCoder.
@@ 4.3.3. Evaluation metric
DepSCode-6.7B 30.0% 23.0% ($\downarrow 23.3\%$) 13.0% 9.0% ($\downarrow 30.8\%$)
DepSCode-33B 31.0% 25.0% ($\downarrow 19.4\%$) 15.5% 11.5%( $\downarrow 25.8\%$)
@@
5. Evaluation Results
Table 3. Static analysis based project context acquisition approaches: model performance under evolving-aware (HumanEvo) and evolving-ignored (EI) settings.
@@ 5. Evaluation Results
5.1. RQ1: How effective is HumanEvo in benchmarking the performance of LLMs in repository-level code generation task?
5.1.1. Overall performance
Table 2 and Table 3 present the performance of seven mainstream LLMs under two different experimental settings: Evolving-Ignored (where the context source is the latest version of the project repository) and Evolving-Aware (HumanEvo, where the context source is the project repository before the target function was committed). Table 2 shows the performance of retrieval based project context acquisition approach, and Table 3 shows the performance of static analysis based project context acquisition approaches.
All the experimental results shown in Table 2 and Table 3 show a significant decrease in performance under the evolving-aware setting compared to the evolving-ignored setting. This suggests that in the evolving-ignored setting, the context source provided to the model contains too much future context, leading to an unrealistically inflated performance of the LLMs.
5.1.2. Break down analysis
Table 4. Break down analysis. EA and EI stand for evolving-aware and evolving-ignored, respectively. ✓means success,
@@ 5.1.2. Break down analysis
We further analyze the generation results of LLMs to investigate whether both issues described in Figure 1 have certain impact on the final results. Table 4 illustrates the overlap of successful task instances between evolving-aware setting and evolving-ignored setting. Since the results from retrieval-based project context acquisition approach and static analysis based project context acquisition approaches are similar, subsequent analysis primarily focuses on the experimental results of the retrieval-based approach. As introduced above, the future context leakage issue will lead to inflated performance, while the useful context leakage issues issue will lead to underestimated performance. However, the results in Table 2 reflect that models’ overall performance only exhibits an inflated trend, driving us to delve deeper into whether the situation of useful context leakage has occurred. We investigate the overlap between task instances that LLMs successfully completed under evolving-ignored and evolving-aware settings.
In Table 4, EA stands for the evolving-aware setting and EI stands for the evolving-ignored setting. A ✓means successful, while a $\times$ means failure. So, EA✓EI $\times$ refers to the number of instances that pass under the evolving-aware setting but fail under the evolving-ignored setting, which is typically caused by useful context missing. EA $\times$ EI✓represents instances that fail under EA but pass under EI, which is typically caused by future context leakage. EA✓EI✓denotes the number of instances that pass under both settings. From the experimental results, it is evident that except for GPT-4 in HumanEvo-Python, which has a value of 0 for EA✓EI $\times$ situation, all other LLMs exhibit deviations from reality caused by the absence of necessary context in a certain quantity, showing that the issue of useful context missing have impact on evolving-ignored setting. However, due to the significantly higher number in EA $\times$ EI✓, the primary influencing factor remains the future context leakage, leading to exaggerated performance of LLMs.
Figure 4. LLM’s performance when using different project version as context source.
Table 5. Model performance in generating functions with deeper dependency levels.
@@
5.1.3. Dependency level
Functions with different dependency levels vary in their reliance on project context, and the evolving-ignored setting directly affects project context. Therefore, in this experiment, We examine the performance differences of LLMs in generating functions with deeper dependency levels, focusing on non-standalone functions in HumanEvo, which encompass intra-class and inter-class functions, across two settings: evolving-ignored and evolving-aware.
@@ 5.1.3. Dependency level
As shown in Table 5, under the evolving-aware setting, the accuracy of LLM in generating both intra-class and inter-class functions has noticeably declined. The decline is more pronounced for inter-class functions, as these functions typically require cross-file context, which is often more scattered. As the project evolves, its various parts tend to become more comprehensive. Therefore, under the evolving-ignored setting, the models can reference an increasing amount of context. Consequently, when the evolving-aware setting loses access to these contexts, the model’s performance significantly deteriorates.
RQ1 Summary: Experiments show an inflated trend in LLM’s performance under the Evolving-ignored setting. Upon further investigation, we discovered that both the presence of too much extraneous context and the disappearance of context that should exist have significantly affected the model’s performance, resulting in the model’s performance deviating from reality.
5.2. RQ2: How does the code generation performance change as repository evolves?
To further illustrate the impact of the evolving-ignored setting on the performance of LLMs as projects evolve, we conduct more in-depth experimental exploration. Typically, a project’s version number consists of three digits in the format like “0.0.0” where the first digit represents the major version, the second digit represents the minor version, and the last digit represents the patch version. To simulate the evolution process of a project, in Figure 4, we start with evo0 (evolving-aware setting) and increment the minor version number by one each time. If there are no more minor versions under that major version, the major version is incremented until the latest version of the project is reached. We do not update the patch version because patch versions generally entail minimal changes to the project.
As shown in Figure 4, the performance of LLMs deviates from reality as the project evolves. However, the artificial improvement in model performance induced by the evolving-ignored setting does not exhibit a linear increase but rather a curve indicating an upward trend. Updates in the project sometimes introduce irrelevant project contexts, which can act as noise and impact the performance of LLMs. Additionally, project updates may result in the removal or modification of originally useful contexts, leading to the direct loss of context that aids in model objective generation, thereby causing a decline in model performance. These factors may be the primary reasons for the occasional slight decrease in model performance when using different versions of the project as context sources.
RQ2 Summary: In the evolving-ignored setting, as the project evolves, the model’s performance gradually deviates from reality, and the model’s overestimation does not increase linearly but rather follows a fluctuating curve.
5.3. RQ3: How different code descriptions in the dataset affect code generation performance?

### v2

@@ III-B Benchmark Construction Pipeline
(1) Project Selection. In order to make HumanEvo more representative, we adopt a strict approach to selecting functions from various open-source projects. For Python, we select high-quality real-world projects from a list of the top 5,000 most downloaded PyPI libraries [^2] https://hugovk.github.io/top-pypi-packages/. Subsequently, we deploy the selected projects and run their testing frameworks to ensure each of them can run correctly and stably. We then proceed to collect PRs from these repositories through the GitHub developer API [^3] https://api.github.com. For Java, as widely recognized projects typically indicate extensive documentation, structured open-source development guidelines, and functional, well-formatted code, our approach involves searching for the top 200 starred projects on GitHub. Then, we deploy every project on the list to select projects that can be successfully deployed and pass the corresponding test suite. Finally, we retain 30 real-world projects, covering over 50 common development domains, which guarantees that the functions selected for evaluation in HumanEvo could represent real-world programming scenarios.
@@
(2) Pull Request Crawling. We crawl a large number of initial PRs from the popular open-source repositories selected in the procedure above. Specifically, for each project, we do not crawl all PRs as we find that early versions of many projects are difficult to deploy successfully. Therefore, for Python, we crawl PRs created after September 2015 (the release time of Python 3.5); for Java, we collect PRs created after September 2014 (the release time of Java 8). Finally, we obtain over 10,000 PRs and meticulously record metadata pertaining to these PRs for subsequent filtering processes.
@@
(3) Attribute-based Filtering. In this step, we preliminarily filter the crawled PRs based on the following three attributes.
@@
  • The PR must introduce new functions. These newly added functions are typically intended to develop the software project by adding new features, improvements, or fulfilling general development objectives, which can be selected as programming task.
@@
(4) Execution-based Filtering. In this step, we perform execution-based filtering for the PRs filtered in the attributes filtering step. Note that, although both the project selection step and the execution-based filtering step have the ”Successful installation” criterion, their purposes are different. The main purpose of requiring successful installation in the project selection step is to quickly filter projects, while in this step, it ensures that all versions associated with different PRs can run successfully, as PRs may correspond to multiple versions of a project.
Firstly, we prepare the execution environment for each candidate task instance in HumanEvo. For each task instance, we first determine the version of the project corresponding to the PR, then manually identify the required dependency environment for running the project at that version. For example, for a Python task instance, we confirm the required Python version and the types and versions of third-party libraries it depends on. After obtaining this dependency environment information, we create a virtual execution environment for each version of the project. Task instances that can not install the dependency environment properly are filtered out during the virtual execution environment setup process. After creating the execution environment, we execute the project’s testing framework in the corresponding environment.
@@
(5) Function Selection. In this step, we manually select functions from the PRs that have passed the filtering process above to serve as programming tasks. After completing the above filtering, we confirm that the remaining PRs meet our requirements. Therefore, we select newly added functions from these PRs as programming tasks. During the selection process, to avoid choosing basic functions, we filter out initialization functions, constructors, and destructors, to ensure the quality of the selected programming tasks. In addition, we ensure that the selected target functions are covered by the project’s test framework through manual review . Finally, we acquire a total of 400 programming tasks.
(6) Human Annotation. In this step, we recruit six annotators with more than five years of programming experience to manually rewrite two different styles of docstrings for the selected functions and label each function with a dependency level. We rewrite the docstring for each instance in HumanEvo to ensure the quality of the benchmark dataset and mitigate data leakage issues as much as possible. Specifically, in the process of data construction, we find that the docstring styles of actual projects can be divided into two categories: the detail and the brief. Docstrings are used as input for models, so variations in docstring style may significantly influence model evaluation. Aiming to determine whether the evolution-ignored setting leads to unreal LLMs’ performance evaluations across different input styles, we provide two types of the docstring for each task instance. The detailed docstring describes the function of the target function in detail, as well as the types and meanings of its inputs and outputs. In contrast, a brief docstring only briefly describes what the target function does. After the annotators complete writing the docstrings, we conduct a cross-validation review, requiring each task instance to be verified and confirmed by two developers to ensure accuracy. If the annotators disagree with a docstring written by the previous annotator during the review process, they will discuss the issue together until all annotators reach an agreement.
@@
III-C Benchmark Characteristics
III-C1 Evolution-aware Benchmark
@@
HumanEvo is the pioneering repository-level code generation benchmark to introduce the evolution-aware concept. We highlight a significant oversight in prior research: the developmental trajectory of a project is an evolution process, altering the project context available to programmers as the project advances through various stages. Previous studies have overlooked this temporal dimension, resulting in LLMs being fed with code that had not yet been implemented in the project when certain functions were developed. Consequently, it also misses some context that should have been present but might have been deleted or updated during the project’s evolution. This undoubtedly leads to the model facing an unrealistic development scenario. As a result, the performance of LLMs in real-world software development tasks remains unknown. Therefore, we are committed to enhancing HumanEvo to serve as a benchmark that better simulates real-world software code development scenarios, aiming to more accurately reflect the true performance of LLMs applied to pragmatic development tasks.
III-C2 Strict Data Filtering Process
@@
We ensure the quality of the selected function from several aspects. First, we ensure that the selected projects are of high quality because high-quality projects generally have excellent maintenance, comprehensive testing frameworks, and good coding styles. For Python, we choose the projects with the highest download count in PyPI. For Java, we select projects from the top 200 projects on GitHub based on the number of stars, with an average of around 24k stars. Previous benchmarks have ensured the functionality of target functions by writing unit tests for each function separately. However, we believe that in a project with complex dependencies, adding a new function may have certain impacts on other parts of the project. Relying solely on unit tests for the target function is insufficient to ensure that the project can still function properly and pass the project’s test framework after inserting the function. Therefore, in this work, we directly run the project’s test framework to ensure the correctness of the target function’s functionality while ensuring that inserting the generated function does not have a negative impact on other parts of the project.
In addition, we also verify that our selected functions are covered by the project’s test framework through execution-based validation. First, we decompose the code diff in the PR corresponding to the target function into a source code patch and a test code patch, and directly apply the test code patch. Then, we validate whether the target function is covered by the project’s test framework by executing the project’s test framework twice. In the first execution, we run the project’s test framework without applying the source code patch. In the second execution, we apply the source code patch and then run the test framework. If there is a test function in the project’s test framework that targets the target function, and the execution of this test function requires calling the target function, the first execution (based on execution) will fail because the target function call is missing. In the second execution, after applying the source code patch, the test framework will correctly call the target function and run as expected. Therefore, if the test framework covers the target function, the results of the two executions should show a “fail-to-pass” transition. We perform execution-based validation for all selected functions to ensure that the HumanEvo test suite is reliable.
TABLE I: Dependency level in HumanEvo.
@@
Standalone 45 (22.5%) 26 (13.0%)
Intra-class 60 (30.0%) 61 (30.5%)
Inter-class 95 (47.5%) 113 (56.5%)
III-C3 Human Annotation
@@
Furthermore, to facilitate the assessment of LLMs’ capability in generating functions with varying levels of dependency, we categorize the selected functions based on their dependency levels, including standalone, intra-class, and inter-class. Standalone functions can be implemented without relying on other functions within the project. Intra-class functions require dependencies on other functions within the same class. Inter-class functions imply more complex dependencies, with most of the dependent functions being located in other files within the project. Manual categorization of dependency levels aids in analyzing the performance of LLMs in generating functions of varying complexity, facilitating researchers in improving the application of LLMs in practical development scenarios based on different contexts. Moreover, as indicated in Table I, the majority of our benchmark consists of functions with deep dependency levels. This suggests that our benchmark presents a sufficiently challenging scenario.
Note that we use PRs to construct HumanEvo primarily for the following three reasons. First, since prompting LLMs to generate repository-level functions can be viewed as an incremental development process, and PRs for adding new features exactly mirror the incremental development process in the real world, we choose to instruct LLMs generate the newly added functions in PRs to simulate real-world development scenarios. Second, the quality of functions in PRs is guaranteed. Before merging the newly added code in PRs into the project, repository administrators conduct a review on it, which allows us to select high-quality task instances directly. Third, the base commit in PRs helps us roll back the project to the state before the target function was committed. Finally, by processing PRs appropriately, we can further determine whether the target functions are covered by the project’s testing framework.
IV Experimental Design
@@
IV-A Studied LLMs
We select the mainstream LLMs (both open-source and closed-source ones) that have been widely used in recent code generation work [68, 13, 64, 6]. For open-source LLMs, we select the CodeLlama series [8] (including CodeLlama-7B, CodeLlama-13B, and CodeLlama-34B) and the DeepSeekCoder series [10] (including DeepSeekCoder-6.7B and DeepSeekCoder-33B). The CodeLlama and DeepSeekCoder series are popular LLMs that perform well in code generation. For closed-source LLMs, we choose the commonly used commercial models: GPT-3.5-Turbo [74] and GPT-4 [16], with GPT-4 demonstrating excellent performance in previous benchmarks.
IV-B Studied Context Acquisition Strategies
Due to the large size of modern software repositories, it is not feasible to feed the entire repository to the model. Thus, previous works mainly obtain project context that might be beneficial for generating target functions through two strategies: retrieval-based and static analysis-based methods. In the following experiments, we cover both of these technical routes to demonstrate that the evolution-ignored situation leads to the deviation of the LLMs’ performance from reality.
Retrieval-based Context Acquisition. We adopt the standard retrieval-augmented-generation pipeline for repository-level code generation [6, 41, 42, 43]. In the evolution-aware setting, we roll back the repository to the state before the target code was committed for each task instance. On the contrary, in the evolution-ignored settings, we directly utilize the latest project version as the context source. Then, we construct a corresponding context retrieval database by partitioning the source code files from the repository into a collection of code snippets. Subsequently, we utilize the docstring of the target code as the query for retrieval and compute the Jaccard similarity between the query and all code snippets in the context retrieval database. After sorting all the code snippets by similarity score, we select project contexts with higher similarity scores as project context for target function generation.
Static Analysis-based Context Acquisition. Similar to the static analysis-based repository-level code generation work [5], we obtain code files in the repository that might help generate the target code through static analysis. First, we roll back the repository to the corresponding state according to evolution-ignored or evolution-aware settings. Then, we extract the source code from these files as project context to feed to the model. There are four types of context sources: local file (the file where the target code resides), import file (files imported by the local file in the project), sibling file (files located in the same folder as the local file), and similar file (files in the project with names similar to the local file). We utilize four prompting strategies from  [5]: local – extracting context solely from the local file; local+import – extracting code from both local and import files; local+sibling - extracting code from both local and sibling files; local+sim - extracting code from both local and similar files.
IV-C Research Questions
Our experiments intend to answer the following research questions (RQs):
@@ IV-C Research Questions
    We conduct experiments under two different settings, evolution-aware and evolution-ignored, to explore the differences in model performance between these two settings and reveal the actual performance of the models in real development scenarios.
@@
    We treat multiple versions of project as context sources to explore how the model’s performance change as the project evolves.
@@
IV-D Experimental Settings
We assess the functionality correctness of code generated by LLMs with execution-based evaluation and utilize the commonly used metric Pass@1 for evaluation. We set the generation temperature of LLMs to 0.8, top-p to 0.95, and the context window length to 4096. Note that, to mitigate issues stemming from the randomness of model generation, the experimental results presented in this paper are obtained by conducting three repeated experiments and averaging the results. The experiment was conducted on a server running Ubuntu 18.04.6 LTS and equipped with 128 Intel(R) Xeon(R) Platinum 8336C @ 2.30GHzE CPUs, and 8 NVIDIA A800 with 80GB memory.
TABLE II: Retrieval-based context acquisition approach: model performance under evolution-aware (HumanEvo) and evolution-ignored (EI) settings.
@@ IV-D Experimental Settings
DeepSeekCoder-6.7B 30.0% 23.0% ($\downarrow 23.3\%$) 13.0% 9.0% ($\downarrow 30.8\%$)
DeepSeekCoder-33B 31.0% 25.0% ($\downarrow 19.4\%$) 15.5% 11.5%( $\downarrow 25.8\%$)
@@
V Evaluation Results
TABLE III: Static analysis based project context acquisition approaches: model performance under evolution-aware (HumanEvo) and evolution-ignored (EI) settings.
@@ V Evaluation Results
V-A RQ1: How effective is HumanEvo in benchmarking the performance of LLMs in repository-level code generation task?
Overall Performance. Table II and Table III present the performance of seven mainstream LLMs under two different experimental settings: Evolution-Ignored (where the context source is the latest version of the project) and Evolution-Aware (HumanEvo, where the context source is the project before the target function was committed). Table II shows the performance of retrieval based project context acquisition approach, and Table III shows the performance of static analysis based project context acquisition approaches. All the experimental results in Table II and Table III show a significant decrease in performance under the evolution-aware setting compared to the evolution-ignored setting. This suggests that ignoring the evolution of the project would lead to inflated performance of LLMs, failing to accurately reflect their real code generation performance.
TABLE IV: Break down analysis. EA and EI stand for evolution-aware and evolution-ignored, respectively. ✓means success,
@@
TABLE V: Model performance in generating functions with deeper dependency levels.
@@
Figure 5: LLM’s performance when using different project versions as context sources.
Break Down Analysis. We further analyze the generation results to investigate whether evolution-ignored setting always leads to over-estimated performance of LLMs. Table IV illustrates the overlap of successful task instances between evolution-aware setting and evolution-ignored setting. Since the results from retrieval-based project context acquisition approach and static analysis based project context acquisition approaches are similar, subsequent analysis primarily focuses on the experimental results of the retrieval-based approach. As introduced above, evolution-ignored setting would raise issues such as the future context leakage issue , which would lead to inflated performance. Meanwhile, evolution-ignored setting can raise issues such as the useful context leakage issue , which would lead to underestimated performance. However, the results in Table II reflect that models’ overall performance only exhibits an inflated trend, driving us to delve deeper into whether issues such as useful context leakage has occurred. We investigate the overlap between task instances that LLMs successfully completed under evolution-ignored and evolution-aware settings.
In Table IV, EA stands for the evolution-aware setting and EI stands for the evolution-ignored setting. A ✓means successful, while a $\times$ means failure. So, EA✓EI $\times$ refers to the number of instances that pass under the evolution-aware setting but fail under the evolution-ignored setting. EA $\times$ EI✓represents instances that fail under EA but pass under EI. EA✓EI✓denotes the number of instances that pass under both settings. From the experimental results, it is evident that except for GPT-4 in HumanEvo-Python, which has a value of 0 for EA✓EI $\times$ situation, all other LLMs exhibit deviations from reality caused by issues such as the useful context missing issue in a certain quantity, showing that the issues , which are raised by the evolution-ignored setting and would lead to underestimated performance, have impact on evolution-ignored setting. However, due to the significantly higher number in EA $\times$ EI✓, the primary influencing factor remains the future context leakage, leading to exaggerated performance of LLMs.
@@
Dependency Level. Functions with different dependency levels vary in their reliance on project context, and the evolution-ignored setting directly affects project context. Therefore, in this experiment, We examine the performance differences of LLMs in generating functions with deeper dependency levels, focusing on non-standalone functions in HumanEvo, which encompass intra-class and inter-class functions, across two settings: evolution-ignored and evolution-aware. As shown in Table V, under the evolution-aware setting, the accuracy of LLM in generating both intra-class and inter-class functions has noticeably declined. The decline is more pronounced for inter-class functions, as these functions typically require cross-file context, which is often more scattered. As the project evolves, its various parts tend to become more comprehensive. Therefore, under the evolution-ignored setting, the models can reference an increasing amount of context. Consequently, when the evolution-aware setting loses access to these contexts, the model’s performance significantly deteriorates.
RQ1 Summary: Experiments show an inflated trend in LLM’s performance under the evolution-ignored setting.
V-B RQ2: How does the code generation performance change as repository evolves?
To further demonstrate the impact of evolution-ignored setting, we treat different versions of the project as context sources to explore how the performance of LLMs change over time. We start from the version right before the target code was committed and incrementally increase the project versions up to the latest version. Typically, a project’s version number consists of two digits in the format “v1.2”, where the first digit denotes the major version and the second digit denotes the minor version. In Figure 5, evo0 is the version right before the target code was committed, standing for the evolution-aware setting. For evo1 to evo10, the version numbers are incrementally increased. For example, if evo0 corresponds to the “v1.2” version of the project, then evo1 will be “v1.3”.
As shown in Figure 5, we find that the performance of LLMs deviates from the actual performance in the evo0 (evolution-aware) setting, with this deviation increasing as the project evolves. All studied LLMs, regardless of being open-source or closed-source and irrespective of parameter size, consistently exhibit inflated performance and follow a similar trend in performance changes. This result further confirms that ignoring project evolution can lead to unrealistic evaluation of LLMs.
RQ2 Summary: In the evolution-ignored setting, as the project evolves, the performance of LLMs deviates from actual performance, with this deviation increasing over time.
V-C RQ3: How different code descriptions in the dataset affect code generation performance?